OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) less than or equal to 25? A. x is greater than or equal to -3 B. x is less than or equal to -3 C. -3 is less than or equal to x 4. A number line with a closed circle on negative 3 and shading to the right.

OpenStudy (anonymous):

7-9x-(x+12)=<25 -7 -7 -9x-(x+12)=<18 -9x-x-12=<18 +12 +12 -10x=<30 divide both sides by -10 and switch the sign x>=-3

OpenStudy (anonymous):

Oh ok that makes sense, I tried solving it but i distributed first.

OpenStudy (anonymous):

Oh and that would mean its B. because it's the only one that doesnt represent it.